How do you change 4x-2y=10 into slope-intercept form? I have no idea what to do. Thanks for your time.
Answer by nerdybill(7384)

How do you change 4x-2y=10 into slope-intercept form?
.
The "slope-intercept" form of a line is:
y = mx + b
all you need to do, is to rearrange the terms of the given equation to make it look like the form above...
You started with:
4x-2y=10
begin by subtracting 4x from both sides:
-4x+4x-2y = -4x+10
-2y = -4x+10
multiply both sides by -1:
2y = 4x-10
divide both sides by 2:
y = 2x-5 (this is what they're looking for)